If S1=0.7 and S2=2.1, in a geometric series, determine the sum of the first 12 terms In the series
Answer by Boreal(15235)

s=a(1-r^n)/(1-r)
r=3 n=12
s=0.7(1-3^12)/-2
=0.7(265720)=186,004
0.7
2.1
6.3
18.9
56.7
170.1
510.3
1530.9
4592.7
13778.1
41334.3
124002.9